OpenStudy (anonymous):

Solve the logarithmic equation for x: log base 2(x+1)+log base 2(x-1)=3

OpenStudy (nowhereman):

Apply exponentiation!

OpenStudy (anonymous):

.........huh? :P

OpenStudy (nowhereman):

If you know the logarithm you should know the exponential.

OpenStudy (anonymous):

I'm lost........ Gah why does math have to be so hard?

OpenStudy (nowhereman):

It's not, if you follow the definitions. How did you define log?

OpenStudy (anonymous):

What do you mean? I didn't define anything... I don't think...

OpenStudy (nowhereman):

I mean in class, or where you learned about the logarithm. Or how do you want to solve any problem containing it, if you don't know what it is?

OpenStudy (anonymous):

I learned this like 6 months ago, so I really don't remember. This is a take home test... I just need to solve for x, I think.

OpenStudy (nowhereman):

Then you should look it up, in your notes actually, but wikipedia is often good too. What you need here is \[b^{\log_b x} = x\] because logarithm is the inverse function of the exponential.

OpenStudy (anonymous):

Is b the base?

OpenStudy (nowhereman):

Indeed

OpenStudy (anonymous):

So it would be \[2^\log2(3)=3\]?

OpenStudy (watchmath):

hi nowhere man, we need to combine the logs first before applying exponentiation. Also in US usually they don't teach exponentiation explicitly in that way.

OpenStudy (anonymous):

That's what I thought, cause I didn't remember anything like that.... Help?

OpenStudy (watchmath):

Do you remember this property? \(log_b A+\log_b B= \log_b(AB)?\)

OpenStudy (anonymous):

I think so...

OpenStudy (watchmath):

Ok, what do you get if you apply that property to the left hand side of your equation?

OpenStudy (anonymous):

\[\log_{2} 2x\]I think. Because the +1 and the -1 would cancel out.

OpenStudy (watchmath):

But you multiply the inside instead of adding them.

OpenStudy (anonymous):

Why would I multiply them? It says I'm adding the two logs...

OpenStudy (anonymous):

Answer x = 3; log(3+1) to the base 2 = 2 and log(3-1) to the base 2 = 1

OpenStudy (watchmath):

look at again \(log_b A +\log_b B=\log_b(AB)\) In there you multiply A and B

OpenStudy (anonymous):

Ohhhhh okay. So it would be\[\log_{2} x^2+1\]

OpenStudy (watchmath):

Check it again what do you get if you do (x+1)(x-1) (foil it :) )

OpenStudy (anonymous):

Formula log(A) to the base 2 + log(B) to the base 2 = log(A B) to the base 2; Steps: 1. log (x+1)(x-1) to the base 2 = 3 2. removing log, 2 raised to 3 = (x+1) (x-1) 3. 8 = x^2 - 1 4. 9 = x^2 5. taking square root on both sides, 3 = x

OpenStudy (anonymous):

\[\log_{2} x^2-1\]

OpenStudy (anonymous):

that would equal 3, right?

OpenStudy (watchmath):

great! So now we have \(\log_2(x^2-1)=3\) Do you remember how to change from log equation into exponential equation?

OpenStudy (anonymous):

I think if I make the three on the other side a log as well I could do it, but that doesn't seem right...

OpenStudy (watchmath):

Remember this \[\log_bx=y\iff b^y=x\]

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

OHH I REMEMBER THAT! :D

OpenStudy (anonymous):

So it would be \[2^3=x^2-1\]

OpenStudy (anonymous):

Yep. Now solve for x :)

OpenStudy (anonymous):

\[x^2=9\]so x=3 :D

OpenStudy (anonymous):

AHHH YOU GUYS ARE AMAZING<3

OpenStudy (anonymous):

x could be -3 as well. Since we are dealing with logarithms, we ignore -3

OpenStudy (anonymous):

Oh yeah, I forgot about that. Negatives don't work with logs.

OpenStudy (watchmath):

To bee a little pedantic from \(x^2=9\) we have \(x=\pm 3\). But when we plug in back x=-3 into the equation we are in trouble because we can't compute log of negative number. SO x=3 is the only solution.

OpenStudy (anonymous):

YAYAYAYYAYAYAYAYYAYAY!<3 Thanks so much guys!

OpenStudy (anonymous):

Any time!